OpenStudy (anonymous):

Find the general solution of the given system of equations: x' = (2, -1; 3,-2)x + (e^t; t) (where ; indicates a new line of the matrix/vector)

OpenStudy (anonymous):

\[x'=\left[\begin{matrix}2 & -1 \\ 3 & -2\end{matrix}\right]x+\left(\begin{matrix}e^t \\ t\end{matrix}\right)\]

OpenStudy (anonymous):

Have you found your eigenvalues/vectors yet?

OpenStudy (anonymous):

Yep! They are \[\left(\begin{matrix}1 \\ 1\end{matrix}\right)\] and \[\left(\begin{matrix}1 \\ 3\end{matrix}\right)\]

OpenStudy (anonymous):

We're supposed to be using undetermined coefficients for this problem to find the particular solution though. I have the solution, but I have no idea how they found it. The final solution should be: \[x=c1\left(\begin{matrix}1 \\ 1\end{matrix}\right)e^t+c2\left(\begin{matrix}1 \\ 3\end{matrix}\right)e^{-t}+\frac{ 3 }{ 2 }\left(\begin{matrix}1 \\ 1\end{matrix}\right)te^t-\frac{ 1 }{ 4 }\left(\begin{matrix}1 \\ 3\end{matrix}\right)e^t+\left(\begin{matrix}1 \\ 2\end{matrix}\right)t-\left(\begin{matrix}0 \\ 1\end{matrix}\right)\]

OpenStudy (anonymous):

Okay, I assume you know how the homogeneous solution was found? For the UC set up, first we split up the nonhomogeneous term like so: \[\begin{pmatrix}e^t\\t\end{pmatrix}=\begin{pmatrix}1\\0\end{pmatrix}e^t+\begin{pmatrix}0\\1\end{pmatrix}t\] and for the guess solution, we use \[x_p=Ae^t+Bt+C\] where \(A,B,C\) are unknown vectors.

OpenStudy (anonymous):

Okay. Although, shouldn't there also be a vector with \[te^t\] in it?

OpenStudy (anonymous):

Ah yes, that's right. I forgot about the homogeneous \(e^t\) solution.

OpenStudy (anonymous):

So \(x_p=Ate^t+Be^t+Ct+D\).

OpenStudy (anonymous):

Okay, sounds good so far

OpenStudy (anonymous):

Then you have \[{x_p}'=Ate^t+(A+B)e^t+C\] Substituting into the original system, \[Ate^t+(A+B)e^t+C=\begin{pmatrix}2&-1\\3&-2\end{pmatrix}(Ate^t+Be^t+Ct+D)+\begin{pmatrix}1\\0\end{pmatrix}e^t+\begin{pmatrix}0\\1\end{pmatrix}t\]

OpenStudy (anonymous):

Then I assume I'd just equate like terms? That would give me \[A=\left[\begin{matrix}2 & -1 \\ 3 & -2\end{matrix}\right]A ; A+B=\left[\begin{matrix}2 & -1 \\ 3 & -2\end{matrix}\right]B+\left(\begin{matrix}1 \\ 0\end{matrix}\right); \left(\begin{matrix}0 \\ 0\end{matrix}\right)=\left[\begin{matrix}2 & -1 \\ 3 & -2\end{matrix}\right]C+\left(\begin{matrix}0 \\ 1\end{matrix}\right)\]

OpenStudy (anonymous):

and \[C=\left[\begin{matrix}2 & -1 \\ 3 & -2\end{matrix}\right]D\]

OpenStudy (anonymous):

I'm not really sure where to go from there though

OpenStudy (anonymous):

That first equation is a bit problematic, though. It suggests \(A=\vec{0}\) :/

OpenStudy (anonymous):

Or just that A is an eigenvector of the matrix corresponding to an eigenvalue of 1

OpenStudy (anonymous):

But that wouldn't explain the \(\dfrac{3}{2}\) in the solution, would it?

OpenStudy (anonymous):

Nope haha... dang.

OpenStudy (anonymous):

Here's a slightly more detailed solution... but still doesn't show the work :\ http://www.math.ucsd.edu/~hus003/math20D/HW6.pdf

OpenStudy (anonymous):

On page 10 of that

OpenStudy (anonymous):

Well, given that \(A\) can be a scalar multiple of the first eigenvalue \(\left(A=\begin{pmatrix}k\\k\end{pmatrix}\text{ for real }k\right)\), I suppose you can use that to your advantage to solve for \(B\), then work backwards to find an appropriate \(k\).

OpenStudy (anonymous):

I think k=3 works

OpenStudy (anonymous):

I'm not really sure how to solve for B without finding k first :\

OpenStudy (anonymous):

I would go ahead with \(A=\dbinom11\) and see what the solution ends up being. I don't see any particular reason to choosing \(k=\dfrac{3}{2}\).

OpenStudy (anonymous):

I'll try that. Otherwise I'll probably just go to bed and ask my professor tomorrow haha... I'm exhausted XD